OpenStudy (anonymous):

Two systems of equations are given below. For each system, choose the best description of its solution. If applicable, give the solution. -x-3y=-6 x+3y-6=0

OpenStudy (imtiaz7):

-x-3y= -6 eq. i x+3y-6=0 > x+3y=6 eq ii From eq. i: -3y= -6+x y= -6+x/-3 eq a put the eq. a in eq ii x +3y=6 then x+3(-6+x/-3) =6 x -18+3x/3 =6 3x-18+3x/3=6 6x-18=12 (by cross multiplication) 6(x-3)=12 (common the 6) x-3= 12/6 x-3= 2 x=2+3 x=5 now we got the value of x so put the value of x in eq i or eq ii as your wish; we put it in eq. ii x+3y=6 5+3y=6 3y=6-5 3y=1 y=1/3 Solved.

OpenStudy (anonymous):

thank you